2x+3y-3=0 (1)
2x+4y-4=0 (2)
solve using sustitution method?

Answer by MathPro(14)

Good Evening,
You are asked to solve the following by substitution:
(1) 2x+3y-3=0
(2) 2x+4y-4=0
Rewrite one of the equations in terms of one variable. It doesn’t matter what you pick. I’ll just pick equation (1) and rewrite so the variable y is isolated:
(1) Subtract 2x from both sides: 2x -2x + 3y-3= -2x
(2) Add 3 to both sides: 3y-3+3=-2x+3
(3) Divide both sides by 3: y=1 -2/3x
Now we have an equation for the variable y in terms of x, so lets substitute this into equation (2):
2x +4(1-(2/3)x) – 4 =0
2x + 4 – (8/3)x -4 =0
2x – (8/3)x = 0
-2/3x=0
So x =0,
Lets substitute x =0 back into any equation to solve for y. I’ll pick equation (1).
2(0) + 3y – 3 =0
Add 3 to both sides: 3y -3 + 3=3
Divide both sides by 3: y = 1
So we have x=0 and y=1. Let’s check this answer by substituting for x and y in any equation. I’ll pick equation (2):
2(0) + 4(1) – 4= 0
0=0
The left-hand-side is equal to the right-hand-side, so the answer is correct.
Good Luck!
Selma from Rent-A-Whiz, an online tutoring and homework help website. Just visit www.rentawhiz.com to learn more. Currently, we are offering to solve one math problem for free. Just visit our website, select the homework tab, then click on the math problem submission link and quote coupon code 003 when you submit your problem.